Justine is ice-skating at the Lloyd Center. What is her final velocity if she accelerates at a rate of

2.0 m/s2 for 3.5 seconds?


1
Expert's answer
2022-03-15T10:47:15-0400

"v_f=v_i+at"

Assuming that she starts from the rest,

"v_f=0\frac{m}{s}+2.0\frac{m}{s^2}×3.5\ s=7.0\frac{m}{s}"


Answer: 7.0 m/s